OpenStudy (anonymous):

Simplify. √7 / 2√11 (fraction)

OpenStudy (jhannybean):

\[\frac{\sqrt{7}}{2\sqrt{11}} \cdot \frac{\sqrt{11}}{\sqrt{11}} = \frac{\sqrt{7\cdot 11}}{2\cdot 11} = ~?\]

OpenStudy (anonymous):

77/22

OpenStudy (jhannybean):

almost.

OpenStudy (jhannybean):

\[\frac{\sqrt{77}}{22}\]

OpenStudy (anonymous):

Thank you! =)))

OpenStudy (jhannybean):

No problem.